Suppose a chinook salmon needs to jump a waterfall that is 1.47 m high.

(a) If the fish starts from a distance 1.13 m from the base of the ledge over which the waterfall flows, find the x- and y-components of the initial velocity the salmon would need to just reach the ledge at the top of its trajectory.

v0x = _____ m/s
v0y = _____ m/s


(b) Can the fish make this jump? (Note that a chinook salmon can jump out of the water with a speed of 6.26 m/s.)

Yes

No     
Answer #1

Vy^2 = uy^2 - 2gs

0 = uy^2 - 2gs

uy = sqrt(2gs) = sqrt(2 x 9.8 x 1.47) = 5.37 m/s

Time, t = 5.37/9.8 = 0.548 sec

Vx = 1.13/0.548 = 2.06 m/s

Vx = 2.06 m/s

Vy = 5.37 m/s